### v1

@@ 1 Introduction
The same inverse proximity effect also arises at interfaces between a superconductor and a ferromagnetic insulator [4]. The key distinction in this case is that, due to the insulating nature of the ferromagnet, the penetration of Cooper pairs into the ferromagnetic region is negligible. The formation of triplet pairs occurs at the expense of singlet pairs, thereby weakening the conventional superconducting state  [5, 6, 7] for thin superconducting films with a thickness smaller than $\xi_{S}$ in contact with a ferromagnet, or results in a suppression of the superconducting order parameter (OP) near the interface over a distance on the order of $\xi_{S}$ in thicker films. The generation of triplet Cooper pairs in superconductor/ferromagnet (S/F) heterostructures has revealed a wealth of new and exciting physics [2, 1] and has provided a strong impetus for the development of superconducting spintronics [8, 9, 3], including the realization of superconducting spin valves [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], Josephson $\pi$-junctions [34, 2] and $\varphi_{0}$-junctions [35, 36], superconducting spin currents carried by triplet Cooper pairs [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], magnetization dynamics and magnetization switching induced by the triplet currents [50, 54, 55, 56, 57, 58, 59, 60, 61], giant thermoelectric effects [62, 63, 64, 65, 66] and efficient low-temperature spin caloritronics [3, 67, 68, 69], hybrid collective excitations [70] and many other phenomena.
On the other hand, antiferromagnets (AFs) are now attracting considerable attention as alternatives to ferromagnets for spintronics applications, owing to their robustness against external magnetic field perturbations, the absence of parasitic stray fields, and their ultrafast dynamics [71, 72, 73]. As discussed above, the majority of spintronics application of S/F heterostructures rely on proximity-induced singlet-triplet conversion at the S/F interfaces.
@@
Recently it was found that in spite of the absence of a net magnetization, the checkerboard Néel magnetic order of the AF does induce spin-triplet correlations at S/AF interfaces, which penetrate into the superconductor and into the antiferromagnet (if it is metallic) [80, 81] on the length scale of $\xi_{S}$ and $\xi_{N}$, respectively. Their amplitude flips sign from one lattice site to the next, just like the Néel magnetic order in the AF. These correlations were called Néel triplet Cooper pairs. The Néel triplet correlations exhibit physics that is in many respects similar to the physics of conventional triplet correlations in S/F heterostructures. In particular, they suppress superconductivity [80]. The efficiency of the suppression is of the same order and even a bit stronger than the suppression by the conventional proximity-induced triplet correlations in S/F heterostructures. Also, the Néel triplet correlations provide an indirect exchange interaction in AF/S/AF trilayers [82, 83], thus opening prospects for antiferromagnet-biased superconducting spintronics. Similar to S/F heterostructures [2, 84, 85, 86, 87, 88, 89, 90, 91, 92], in S/AF heterostructures a mesoscopic FFLO state was predicted [93, 94, 81], which experimentally should manifest itself as oscillations of the critical temperature of S/AF bilayers depending on the thickness of the AF layer [93], which seem to be similar to existing experimental observations [95, 96, 97, 98]. However, the physical mechanism of the formation of the mesoscopic FFLO state in S/AF heterostructures [81] is not based on the nonzero macroscopic magnetization and is completely different from S/F heterostructures [99]. By analogy with the S/F heterostructures [100, 101, 102, 103, 104, 105, 106] the Néel triplets result in the anisotropy of the critical temperature of S/AF heterostructures in the presence of spin-orbit coupling (SOC) [107].
At the same time, the Néel triplet correlations exhibit significant differences from conventional triplet correlations induced by the proximity effect in S/F heterostructures. Unlike in S/F heterostuctures, the amplitude of the Néel triplet correlations is highly sensitive to the value of the chemical potential (filling factor) in the material in which they are induced. Deviations from half filling suppress the amplitude of the Néel triplets for a given value of the antiferromagnet exchange field [80].
@@
In this work, we demonstrate that the appearance of the Néel triplet correlations in antiferromagnetic superconductors and S/AF heterostructures is also highly sensitive to the type of the antiferromagnet order. All the physics discussed above was related to the checkerboard (G-type) antiferromagnetic order. In the present study, we extend the analysis to A-type and C-type (layered) antiferromagnets and compare the resulting superconducting proximity effects to those in S/F heterostructures and in S/AF heterostructures with checkerboard antiferromagnets. Our analysis includes both 2D and 3D geometries and considers different orientations of the antiferromagnet relative to the S/AF interface. Unexpectedly, we find that noticeable A-type or C-type Néel triplet correlations are not generated in antiferromagnetic superconductors and S/AF heterostructures even at S/AF interfaces with A-type or C-type antiferromagnets or in antiferromagnetic superconductors with A-type or C-type antiferromagnetic ordering. Instead, the layered antiferromagnets can only produce weak conventional triplet correlations or checkerboard triplet correlations depending on the orientation of the S/AF interface. Moreover, in some physical situations the proximity-induced triplets are not generated at all. Correspondingly, the suppression of the superconducting OP by such types of antiferromagnetism at S/AF interfaces is much weaker than in the S/F heterostructures or S/AF heterostructures with G-type antiferromagnets. We analyze the physical origin of these seemingly counterintuitive results.
@@ 2 Model and generalized two-sublattice formalism
where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})^{T}$ is the vector of Pauli matrices in spin space, $\nu=A,B$ is the sublattice index, $\hat{\psi}_{\bm{i}\sigma}^{\nu\dagger}(\hat{\psi}_{\bm{i}\sigma}^{\nu})$ is the creation (annihilation) operator for an electron with spin $\sigma=\uparrow,\downarrow$ at the sublattice $\nu$ of the unit cell $\bm{i}$ what we will denote as $(\bm{i},\nu)$. $t_{\bm{i}\bm{j}}^{\nu\nu^{\prime}}$ parameterizes the hopping between sites $(\bm{i},\nu)$ and $(\bm{j},\nu^{\prime})$, $\Delta_{\bm{i}}^{\nu}$ accounts for on-site s-wave pairing. For homogeneous antiferromagnetic superconductors it is nonzero at all sites of the system, and for S/AF heterostructures it is only nonzero in the S layer. $\hat{n}_{\bm{i}\sigma}^{\nu}=\hat{\psi}_{\bm{i}\sigma}^{\nu\dagger}\hat{\psi}_{\bm{i}\sigma}^{\nu}$ is the particle number operator at the site $(\bm{i},\nu)$. $h_{\bm{i}}$ is an exchange term induced by localized magnetization at the site $(\bm{i},\nu)$. Again, for homogeneous antiferromagnetic superconductors it is nonzero at all sites of the system, and for S/AF heterostructures it only presents inside the antiferromagnet. Note that the filling factor, often denoted as $\mu_{\bm{i}}$, is also included in this first term (on-site term $\mu_{\bm{i}}=2t_{\bm{i}\bm{i}}^{\nu\nu}$).
The Matsubara Green’s function in the two-sublattice formalism is $\times 8$ 8 matrix in the direct product of spin, particle-hole and sublattice spaces. Introducing the two-sublattice Nambu spinor $\check{\psi}_{\bm{i}}=(\hat{\psi}_{\bm{i},\uparrow}^{A},\hat{\psi}_{\bm{i},\downarrow}^{A},\hat{\psi}_{\bm{i},\uparrow}^{B},\hat{\psi}_{\bm{i},\downarrow}^{B},\hat{\psi}_{\bm{i},\uparrow}^{A\dagger},\hat{\psi}_{\bm{i},\downarrow}^{A\dagger},\hat{\psi}_{\bm{i},\uparrow}^{B\dagger},\hat{\psi}_{\bm{i},\downarrow}^{B\dagger})^{T}$ we define the Green’s function as follows:
@@
where $\langle T_{\tau}...\rangle$ means imaginary time-ordered thermal averaging. For the system described by Hamiltonian (1) the Heisenberg equation of motion for spinor $\check{\psi}_{\bm{i}}$ takes the form:
$\frac{d\check{\psi}_{\bm{i}}}{d\tau}=[\hat{H},\check{\psi}_{\bm{i}}]=\tau_{z}\Bigl{(}\hat{j}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\hat{\bm{h}_{i}}\check{\bm{\sigma}}\Bigr{)}\check{\psi}_{\bm{i}},$ (3)
where $\sigma_{k}$, $\tau_{k}$ and $\rho_{k}$ ($k=x,y,z$) are Pauli matrices in spin, particle-hole and sublattice spaces, respectively. $\hat{\bm{h}}=\bm{h}^{A}(1+\rho_{z})/2+\bm{h}^{B}(1-\rho_{z})/2$, $\check{\bm{\sigma}}=\bm{\sigma}(1+\tau_{z})/2+\bm{\sigma}^{*}(1-\tau_{z})/2$ is the quasiparticle spin operator and $\check{\Delta}_{\bm{i}}=\Delta_{\bm{i}}\tau_{+}+\Delta_{\bm{i}}^{*}\tau_{-}$ with $\tau_{\pm}=(\tau_{x}\pm i\tau_{y})/2$. Here we assume that $\Delta_{\bm{i}}^{A}=\Delta_{\bm{i}}^{B}=\Delta_{\bm{i}}$, that is the order parameter values are equal for $A$ and $B$ sites of a unit cell which is a consequence of equivalence of two sublattices and is confirmed by the results of the numerical solution of the self-consistency equation.
$\hat{j}\check{\psi}_{\bm{i}}=\sum\limits_{\bm{j}}\left(\begin{array}[]{cc}t_{\bm{i}\bm{j}}^{AA}&t_{\bm{i}\bm{j}}^{AB}\\ t_{\bm{i}\bm{j}}^{BA}&t_{\bm{i}\bm{j}}^{BB}\end{array}\right)_{\rho}\check{\psi}_{\bm{j}}.$ (6)
@@
$G_{\bm{i}}^{-1}=\tau_{z}(\hat{j}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\hat{\bm{h}}_{\bm{i}}\check{\bm{\sigma}})+i\omega_{m}.$ (9)
@@ 3.1 Triplet correlations in a homogeneous antiferromagnetic superconductor
$\left[i\omega_{m}\tau_{z}+\tau_{z}\check{\Delta}-\hat{\bm{h}}\bm{\sigma}\tau_{z}\right]\rho_{x}\check{\tilde{G}}(\bm{p})+\left(\begin{array}[]{cc}\xi^{AA}(\bm{p})&\xi^{AB}(\bm{p})\\ \xi^{BA}(\bm{p})&\xi^{BB}(\bm{p})\end{array}\right)_{\rho}\check{\tilde{G}}(\bm{p})=1.$ (22)
@@
where $\xi_{1,2}=(\xi_{D}\pm\xi_{O})$. The dispersion relation of electrons in the superconducting antiferromagnet can be taken from the pole of the Green’s function Eq. (23). In the normal state $\Delta=0$ the electron dispersion takes the form $\varepsilon=-\frac{1}{2}(\xi_{D}\pm\sqrt{\xi_{O}^{2}+h^{2}})$. Thus, physically $-\xi_{1,2}$ are two branches of the electron dispersion at $h=0$.
@@
$F_{x,y,z}={\rm O}[\frac{1}{t}]$ (26)
[…]
@@ 3.2.1 BdG approach
In this section we study the triplet correlations induced at S/AF interfaces, that is in the case when superconductivity and antiferromagnetism are spatially separated. Namely, we consider a S/AF heterostructure infinite along the $y$ and $z$ axes for the 3D case or along the $y$ axis for the 2D case in the simplest tight-binding model on a square lattice with only nonzero nearest-neighbor hopping element $t$. Also we restrict ourselves by the half-filling case $\mu=0$ in order to avoid Friedel oscillations of the correlation functions and absolutely transparent S/AF interface. This spatially inhomogeneous problem is studied numerically and for the numerical calculations we do not use the two-sublattice formalism described above. Then the corresponding Hamiltonian takes the form:
@@
The superconducting order parameter in is calculated self-consistently:
@@
Since the problem is translationally invariant, we choose a box of 2 atoms in the $y$ direction in 2D (of $\times 2$ 2 atoms in the $yz$ plane for 3D) and also consider the full coordinate dependence along the $x$-axis. The spatial dependence of the solution along the $y$ and $z$ directions takes the form:
@@
where $\bm{\eta}$ is the projection of $\bm{i}$ on the chosen translationally invariant box. Then Eqs. (30) and (31) take the form:
@@
$F_{\bm{\eta},\alpha\beta}(\omega_{m})=\int\limits_{BZ}\frac{d(pa)^{D-1}}{\pi^{D-1}}\sum\limits_{n}\left(\frac{u_{n,\alpha}^{\bm{\eta}}(\bm{p})v_{n,\beta}^{\bm{\eta}*}(\bm{p})}{i\omega_{m}-\varepsilon_{n}(\bm{p})}+\frac{u_{n,\beta}^{\bm{\eta}}(\bm{p})v_{n,\alpha}^{\bm{\eta}*}(\bm{p})}{i\omega_{m}+\varepsilon_{n}(\bm{p})}\right),$ (35)
where $D$ is the dimensionality of the problem.

### v2

@@ 1 Introduction
The same inverse proximity effect also arises at interfaces between a superconductor and a ferromagnetic insulator [4]. The key distinction in this case is that, due to the insulating nature of the ferromagnet, the penetration of Cooper pairs into the ferromagnetic region is negligible. The formation of triplet pairs occurs at the expense of singlet pairs, thereby weakening the conventional superconducting state  [5, 6, 7] for thin superconducting films with a thickness smaller than $\xi_{S}$ in contact with a ferromagnet, or results in a suppression of the superconducting order parameter (OP) near the interface over a distance on the order of $\xi_{S}$ in thicker films. The generation of triplet Cooper pairs in superconductor/ferromagnet (S/F) heterostructures has revealed a wealth of new and exciting physics [2, 1] and has provided a strong impetus for the development of superconducting spintronics [8, 9, 3], including the realization of superconducting spin valves [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], Josephson $\pi$-junctions [34, 2] and $\varphi_{0}$-junctions [35, 36, 37], superconducting spin currents carried by triplet Cooper pairs [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], superconducting spin diode effect [55, 56], magnetization dynamics and magnetization switching induced by the triplet currents [51, 57, 58, 59, 60, 61, 62, 63, 64], giant thermoelectric effects [65, 66, 67, 68, 69] and efficient low-temperature spin caloritronics [3, 70, 71, 72], hybrid collective excitations [73] and many other phenomena.
On the other hand, antiferromagnets (AFs) are now attracting considerable attention as alternatives to ferromagnets for spintronics applications, owing to their robustness against external magnetic field perturbations, the absence of parasitic stray fields, and their ultrafast dynamics [74, 75, 76]. As discussed above, the majority of spintronics applications of S/F heterostructures rely on proximity-induced singlet-triplet conversion at the S/F interfaces.
@@
Recently it was found that in spite of the absence of a net magnetization, the checkerboard Néel magnetic order of the AF does induce spin-triplet correlations at S/AF interfaces, which penetrate into the superconductor and into the antiferromagnet (if it is metallic) [83, 84] on the length scale of $\xi_{S}$ and $\xi_{N}$, respectively. Their amplitude flips sign from one lattice site to the next, just like the Néel magnetic order in the AF. These correlations were called Néel triplet Cooper pairs. The Néel triplet correlations exhibit physics that is in many respects similar to the physics of conventional triplet correlations in S/F heterostructures. In particular, they suppress superconductivity [83]. The efficiency of the suppression is of the same order and even a bit stronger than the suppression by the conventional proximity-induced triplet correlations in S/F heterostructures. Also, the Néel triplet correlations provide an indirect exchange interaction in AF/S/AF trilayers [85, 86], thus opening prospects for antiferromagnet-based superconducting spintronics. Similar to S/F heterostructures [2, 87, 88, 89, 90, 91, 92, 93, 94, 95], in S/AF heterostructures a mesoscopic FFLO state was predicted [96, 97, 84], which experimentally should manifest itself as oscillations of the critical temperature of S/AF bilayers depending on the thickness of the AF layer [96], which seem to be similar to existing experimental observations [98, 99, 100, 101]. However, the physical mechanism of the formation of the mesoscopic FFLO state in S/AF heterostructures [84] is not based on the nonzero macroscopic magnetization and is completely different from S/F heterostructures [102]. By analogy with the S/F heterostructures [103, 104, 105, 106, 107, 108, 109] the Néel triplets result in the anisotropy of the critical temperature of S/AF heterostructures in the presence of spin-orbit coupling (SOC) [110].
At the same time, the Néel triplet correlations exhibit significant differences from conventional triplet correlations induced by the proximity effect in S/F heterostructures. Unlike in S/F heterostuctures, the amplitude of the Néel triplet correlations is highly sensitive to the value of the chemical potential (filling factor) in the material in which they are induced. Deviations from half filling suppress the amplitude of the Néel triplets for a given value of the antiferromagnet exchange field [83]. In particular, if the antiferromagnetic exchange field $h$ is small with respect to the chemical potential $\mu$, the Néel triplets are of the order of $h/\mu\ll 1$. In with case they can be disregarded and any triplet correlations, both short-range and long-range, can only be of conventional type and are only produced if the S/AF interface is uncompemsated [81, 111].
@@
In this work, we demonstrate that the appearance of the Néel triplet correlations in antiferromagnetic superconductors and S/AF heterostructures is also highly sensitive to the type of the antiferromagnet order. All the physics discussed above was related to the checkerboard (G-type) antiferromagnetic order. In the present study, we extend the analysis to A-type and C-type (layered) antiferromagnets and compare the resulting superconducting proximity effects to those in S/F heterostructures and in S/AF heterostructures with checkerboard antiferromagnets. Our analysis includes both 2D and 3D geometries and considers different orientations of the antiferromagnet relative to the S/AF interface. Unexpectedly, we find that noticeable A-type or C-type Néel triplet correlations are not generated in antiferromagnetic superconductors and S/AF heterostructures even at S/AF interfaces with A-type or C-type antiferromagnets or in antiferromagnetic superconductors with A-type or C-type antiferromagnetic ordering and $s$-wave superconducting pairing. Instead, the layered antiferromagnets can only produce weak conventional triplet correlations or checkerboard triplet correlations depending on the orientation of the S/AF interface. Moreover, in some physical situations the proximity-induced triplets are not generated at all. Correspondingly, the suppression of the superconducting OP by such types of antiferromagnetism at S/AF interfaces is much weaker than in the S/F heterostructures or S/AF heterostructures with G-type antiferromagnets. We analyze the physical origin of these seemingly counterintuitive results.
@@ 2 Model and generalized two-sublattice formalism
where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})^{T}$ is the vector of Pauli matrices in spin space, $\nu=A,B$ is the sublattice index, $\hat{\psi}_{\bm{i}\sigma}^{\nu\dagger}(\hat{\psi}_{\bm{i}\sigma}^{\nu})$ is the creation (annihilation) operator for an electron with spin $\sigma=\uparrow,\downarrow$ at the sublattice $\nu$ of the unit cell $\bm{i}$ what we will denote as $(\bm{i},\nu)$. $t_{\bm{i}\bm{j}}^{\nu\nu^{\prime}}$ parameterizes the hopping between sites $(\bm{i},\nu)$ and $(\bm{j},\nu^{\prime})$, $\Delta_{\bm{i}}^{\nu}$ accounts for on-site s-wave pairing. For homogeneous antiferromagnetic superconductors it is nonzero at all sites of the system, and for S/AF heterostructures it is only nonzero in the S layer. It is worth mentioning that we only consider antiferromagnetic superconductors with $s$-wave pairing [79] here. At the same time, other types of superconducting pairing can also be realized in antiferromagnetic superconductors. In particular, if one studies a spin-fluctuation-mediated superconductivity in the framework of the Habbard model, the resulting dominating pairing can have a $d$-wave momentum symmetry [114]. In this case study of the singlet-triplet conversion is a separate problem, which is beyond the scope of our work. $\hat{n}_{\bm{i}\sigma}^{\nu}=\hat{\psi}_{\bm{i}\sigma}^{\nu\dagger}\hat{\psi}_{\bm{i}\sigma}^{\nu}$ is the particle number operator at the site $(\bm{i},\nu)$. $h_{\bm{i}}$ is an exchange term induced by localized magnetization at the site $(\bm{i},\nu)$. Again, for homogeneous antiferromagnetic superconductors it is nonzero at all sites of the system, and for S/AF heterostructures it only presents inside the antiferromagnet. Note that the filling factor, often denoted as $\mu_{\bm{i}}$, is also included in this first term (on-site term $\mu_{\bm{i}}=2t_{\bm{i}\bm{i}}^{\nu\nu}$).
The Matsubara Green’s function in the two-sublattice formalism is $8\times 8$ matrix in the direct product of spin, particle-hole and sublattice spaces. Introducing the two-sublattice Nambu spinor $\check{\psi}_{\bm{i}}=(\hat{\psi}_{\bm{i},\uparrow}^{A},\hat{\psi}_{\bm{i},\downarrow}^{A},\hat{\psi}_{\bm{i},\uparrow}^{B},\hat{\psi}_{\bm{i},\downarrow}^{B},\hat{\psi}_{\bm{i},\uparrow}^{A\dagger},\hat{\psi}_{\bm{i},\downarrow}^{A\dagger},\hat{\psi}_{\bm{i},\uparrow}^{B\dagger},\hat{\psi}_{\bm{i},\downarrow}^{B\dagger})^{T}$ we define the Green’s function as follows:
@@
where $\langle T_{\tau}$$\rangle$ means imaginary time-ordered thermal averaging. For the system described by Hamiltonian (1) the Heisenberg equation of motion for spinor $\check{\psi}_{\bm{i}}$ takes the form:
$\frac{d\check{\psi}_{\bm{i}}}{d\tau}=[\hat{H},\check{\psi}_{\bm{i}}]=\tau_{z}\Bigl({\hat{t}}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\hat{\bm{h}_{i}}\check{\bm{\sigma}}\Bigr)\check{\psi}_{\bm{i}},$ (3)
where $\sigma_{k}$, $\tau_{k}$ and $\rho_{k}$ ($k=x,y,z$) are Pauli matrices in spin, particle-hole and sublattice spaces, respectively. $\hat{\bm{h}}=\bm{h}^{A}(1+\rho_{z})/2+\bm{h}^{B}(1-\rho_{z})/2$, $\check{\bm{\sigma}}=\bm{\sigma}(1+\tau_{z})/2+\bm{\sigma}^{*}(1-\tau_{z})/2$ is the quasiparticle spin operator and $\check{\Delta}_{\bm{i}}=\Delta_{\bm{i}}\tau_{+}+\Delta_{\bm{i}}^{*}\tau_{-}$ with $\tau_{\pm}=(\tau_{x}\pm i\tau_{y})/2$. Here we assume that $\Delta_{\bm{i}}^{A}=\Delta_{\bm{i}}^{B}=\Delta_{\bm{i}}$, that is the order parameter values are equal for $A$ and $B$ sites of a unit cell which is a consequence of equivalence of two sublattices and is confirmed by the results of the numerical solution of the self-consistency equation. The operator $\hat{t}$ is defined as
$\hat{t}\check{\psi}_{\bm{i}}=\sum\limits_{\bm{j}}\left(\begin{array}[]{cc}t_{\bm{i}\bm{j}}^{AA}&t_{\bm{i}\bm{j}}^{AB}\\ t_{\bm{i}\bm{j}}^{BA}&t_{\bm{i}\bm{j}}^{BB}\end{array}\right)_{\rho}\check{\psi}_{\bm{j}}.$ (6)
@@
$G_{\bm{i}}^{-1}=\tau_{z}(\hat{t}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\hat{\bm{h}}_{\bm{i}}\check{\bm{\sigma}})+i\omega_{m}.$ (9)
@@ 3.1 Triplet correlations in a homogeneous antiferromagnetic superconductor
$\left[i\omega_{m}\tau_{z}+\tau_{z}\check{\Delta}-\hat{\bm{h}}\bm{\sigma}\tau_{z}\right]\check{\tilde{G}}(\bm{p})+\left(\begin{array}[]{cc}\xi^{AA}(\bm{p})&\xi^{AB}(\bm{p})\\ \xi^{BA}(\bm{p})&\xi^{BB}(\bm{p})\end{array}\right)_{\rho}\check{\tilde{G}}(\bm{p})=1.$ (22)
@@
where $\xi_{1,2}=(\xi_{D}\pm\xi_{O})$. The dispersion relation of electrons in the superconducting antiferromagnet can be taken from the pole of the Green’s function Eq. (23). In the normal state $\Delta=0$ the electron dispersion takes the form $\varepsilon=-\frac{1}{2}(\xi_{D}\pm\sqrt{\xi_{O}^{2}+h^{2}})$. Thus, physically $\xi_{1,2}$ are two branches of the electron dispersion at $h=0$.
@@
$F_{x,y,z}={\rm O}[\frac{1}{t}]$ . (26)
Up to the leading order with respect to the parameter ${\rm max}(h,\Delta,T)/t$ solutions for the anomalous Green’s functions expressed by Eqs. (25) and (26) coincide with the solution for a conventional superconductor in the absence of an exchange field. This means that in this case the triplet correlations of any type are not produced to the considered accuracy. For the majority of materials and especially for thin-film S/AF heterostructures, which are frequently considered in the framework of the simplified model of the homogeneous antiferromagnetic superconductor [83, 112, 115] the condition ${\rm max}(h,\Delta,T)/t\ll 1$ is fulfilled with good accuracy because $t$ is of the order of eV, while $h$ varies from $1$ up to $100$ meV and $\Delta$ is typically even smaller. Since the triplet correlations are not produced and the singlet-triplet conversion is absent, there is $\it no$ suppression of superconductivity. More precisely, it is small in parameter $h/t]^{2}$.
@@ 3.2.1 BdG approach
In this section we study the triplet correlations induced at S/AF interfaces, that is in the case when superconductivity and antiferromagnetism are spatially separated. Namely, we consider a S/AF heterostructure infinite along the $y$ and $z$ axes for the 3D case or along the $y$ axis for the 2D case in the simplest tight-binding model on a square lattice with only nonzero nearest-neighbor hopping element $t$. Also we restrict ourselves by the half-filling case $\mu=0$ in order to avoid Friedel oscillations of the correlation functions and absolutely transparent S/AF interface. This spatially inhomogeneous problem is studied numerically and for the numerical calculations we do not use the two-sublattice formalism described above. Instead, we choose a box of 2 atoms along the $y$-axis direction in 2D (of $2\times 2$ atoms in the $yz$-plane for 3D), then make a Fourier transform along the S/AF interface and consider the full coordinate dependence along the $x$-axis. This approach is fully equivalent to the two-sublattice formalism, but a bit more convenient for numerical calculations. Then the corresponding Hamiltonian takes the form:
@@
The superconducting order parameter in Eqs. (29) is calculated self-consistently:
@@
Since the problem is translationally invariant, we choose a box of 2 atoms in the $y$ direction in 2D (of $2\times 2$ atoms in the $yz$ plane for 3D) and also consider the full coordinate dependence along the $x$-axis. The spatial dependence of the solution along the $y$ and $z$ directions takes the form:
@@
where $\bm{\eta}$ is the projection of $\bm{i}$ on the chosen translationally invariant box. In this case, solution of Eq. (29) is reduced to a problem of finding the eigenvalues and eigenvectors of $2^{D+1}(N_{S}+N_{AF})\times 2^{D+1}(N_{S}+N_{AF})$ matrix (here $D$ is the dimensionality of the problem and $N_{S},N_{AF}$ is the number of sites along $x$-axis for superconducting and magnetic parts of the system). The eigenvalues of this matrix are the energies $\varepsilon_{n}$ and its eigenvectors consist of the amplitudes $u_{n,\sigma}^{\bm{\eta}},v_{n,\sigma}^{\bm{\eta}}$ for all sites $\bm{\eta}$. By numerically diagonalizing the matrix, we find its eigenenergies and eigenvectors. Then Eqs. (30) and (31) take the form:
@@
$F_{\bm{\eta},\alpha\beta}(\omega_{m})=\int\limits_{BZ}\frac{d(pa)^{D-1}}{\pi^{D-1}}\sum\limits_{n}\left(\frac{u_{n,\alpha}^{\bm{\eta}}(\bm{p})v_{n,\beta}^{\bm{\eta}*}(\bm{p})}{i\omega_{m}-\varepsilon_{n}(\bm{p})}+\frac{u_{n,\beta}^{\bm{\eta}}(\bm{p})v_{n,\alpha}^{\bm{\eta}*}(\bm{p})}{i\omega_{m}+\varepsilon_{n}(\bm{p})}\right).$ (35)